OpenStudy (anonymous):

Max can mow a lawn in 45 minutes. Jan takes twice as long to mow the same lawn. If they work together, how long will it take them to mow the lawn?

OpenStudy (mathstudent55):

Look at each worker's rate of work. I prefer to convert to hours. Max does the job in 45 minutes or 0.75 hour His rate is 1/0.75 job/hour Jan does the job in 90 minutes or 1.5 hours. Her rate is 1/1.5 job per hour When you add these rates, you get the total rate, x. \(\dfrac{1}{0.75} + \dfrac{1}{1.5} = \dfrac{1}{x} \) Solve for x.